OpenStudy (anonymous):

Giving Medal! <3 What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7

OpenStudy (_michaeld22):

-98/7 = -14. -8 - 6 = -14. It's A.

OpenStudy (anonymous):

Thanks! I thought so too.